A 24-hour music network plans to add satellite technology to allow its expansion into other major southern markets. The network expects the quarterly revenue to increase by $380,000 from new cable subscription fees. The satellite will require an initial cost of $1 million with quarterly operating and maintenance costs of $340,000. It will have a $196,170 salvage value after 6 years. Calculate the net present worth of this investment at an interest rate of 4% per year, compounded quarterly.
Answer #1

If Annual interest rate is = 4%

Then, Quarterly interest rate is = 4% / 4 = 1%

Total Number of quarter in 6 years = n = 6 *4 = 24

The Net quarterly revenue = Revenue - maintenance cost = 380000 - 340000 = 40000

Therefore,

Net present worth = -P + A (P / A , i , n ) + F ( P / F , i , n )

= - 1000000 + 40000 ( P / A , 1% , 24 ) + 196170 ( P / F , 1% , 24 )

= - 1000000 + 40000 ( 21.2434 ) + 196170 ( 0.7876 )

= - 1000000 + 849736 + 154503.492

= - 1000000 + 1004239.492

= - $4239.492